OpenStudy (anonymous):

convert recurring decimal 6.838383 into a fraction

OpenStudy (anonymous):

Did you learn anything?

OpenStudy (anonymous):

nop not really it was too much and confusing

OpenStudy (anonymous):

plus thats a differrent question

OpenStudy (anonymous):

It's no different than the earlier one if you follow the same steps

OpenStudy (anonymous):

i copied it anyway

OpenStudy (anonymous):

it is cuz it doesnt start with zero

OpenStudy (anonymous):

That's where you utilize your thinking skills

OpenStudy (anonymous):

What grade are you?

OpenStudy (anonymous):

I'll walk you through it. The idea is to get rid of the repeated decimals by multiplying by 10s or 100s

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

suppose x=6.8383

OpenStudy (anonymous):

what's 10x?

OpenStudy (anonymous):

68.838383?

OpenStudy (anonymous):

No. How did you get that?

OpenStudy (anonymous):

dont you multiply it by 10

OpenStudy (anonymous):

you did that on the first one

OpenStudy (anonymous):

Yes, but that's not the same thing. Multiply 6.8383 by 10

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

what do you get?

OpenStudy (anonymous):

i got the answer 68.383

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so we have (1), x=6.8383 (2),10x=68.3838 Right? Subtract and you'll get: 9x=?

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

You realize you still have the decimals. so you try 100

OpenStudy (anonymous):

->100x=683.8383

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Remember the idea is to get rid of the decimals right

OpenStudy (anonymous):

683.8383 or 68.3838

OpenStudy (anonymous):

x=6.8383 100x=683.8383

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Now subtract and you'll get: 99x=?

OpenStudy (anonymous):

You subtract like terms: so 100x-x=99x 683.8383-6.8383=?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

I want you give the answer

OpenStudy (anonymous):

677

OpenStudy (anonymous):

okay. So this means, 99x=677 therefore x=?

OpenStudy (anonymous):

6.8383?

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

i dont know haha i swear i cant get it like i dont know what your question is

OpenStudy (anonymous):

99x=677 so what is x?

OpenStudy (anonymous):

99? 677?

OpenStudy (anonymous):

What grade are you again?

OpenStudy (anonymous):

why would you ask

OpenStudy (anonymous):

So I know what level math you are and how best to help you

OpenStudy (anonymous):

are you a teacher?

OpenStudy (anonymous):

10th

OpenStudy (anonymous):

No I'm not a teacher by profession. But why would you lie?

OpenStudy (anonymous):

lie about what?

OpenStudy (anonymous):

well, the answer you look for is 677/99

OpenStudy (anonymous):

677/99=6.8383..